Barbour Corporation, located in Buffalo, New York, is a retailer of high-tech products and is known for its excellent quality and innovation. Recently, the firm conducted a relevant cost analysis of one of its product lines that has only two products, T-1 and T-2. The sales for T-2 are decreasing and the purchase costs are increasing. The firm might drop T-2 and sell only T-1.

Barbour allocates fixed costs to products on the basis of sales revenue. When the president of Barbour saw the income statements (see below), he agreed that T-2 should be dropped. If T-2 is dropped, sales of T-1 are expected to increase by 10% next year, but the firm’s cost structure will remain the same.

T-1 T-2
Sales $ 235,000 $ 288,000
Variable costs:
Cost of goods sold 77,000 144,000
Selling & administrative 17,000 57,000
Contribution margin $ 141,000 $ 87,000
Fixed expenses:
Fixed corporate costs 67,000 82,000
Fixed selling and administrative 19,000 28,000
Total fixed expenses $ 86,000 $ 110,000
Operating income $ 55,000 $ (23,000 )

Required:

1. Find the expected change in annual operating income by dropping T-2 and selling only T-1.

2. By what percentage would sales from T-1 have to increase in order to make up the financial loss from dropping T-2? (Enter your answer as a percentage rounded to 2 decimal places (i.e. 0.1234 should be entered as 12.34).)

3. What is the required percentage increase in sales from T-1 to compensate for lost margin from T-2, if total fixed costs can be reduced by $50,000? (Enter your answer as a percentage rounded to 2 decimal places (i.e. 0.1234 should be entered as 12.34).)

1.2.Required % increase in sales of T-1%3.Required % increase in sales from T-1%
Answer #1

1.Expected Change in annual income = Additional Contribution margin from T-1 less contribution margin from T-2

= 141,000*10% - 87,000

= $(72,900)

i.e. income will reduce by $72,900

2.% increase required = 87,000/141,000 = 61.70%

3.% increase required = (Lost margin – fixed costs saved)/Margin from T-1

= (87,000-50,000)/141,000

= 26.24%